# biblio.bib

### v1

@@ Appendix C Dataset details
In this section, we present the details of the training dataset used to train the model in Section 4.
@@
We create the dataset by generating random circuits consisting of the gates $\left\{\mathrm{h},\mathrm{cx},\mathrm{ccx},\mathrm{swap},\mathrm{rx},\mathrm{ry},\mathrm{rz},\mathrm{cp}\right\}$ with CUDA-Q [cudaq] (see details on the latter below). For this, we first sample a subset of the gates from the previous set, and then append these gates uniformly to an empty circuit. We do this randomly from 4 to 32 gates and for 3 to 5 qubits. Once the gate type is fixed, we sample the continuous parameters, of the parameterized gates, following a uniform distribution. After a circuit is sampled, we evaluate the implemented unitary and store the circuit-unitary pair. Additionally, we parse the sampled gate subset into a text prompt and store it alongside the circuits.
We sample random circuits until our dataset consists of 11 million (M) unique circuit ansätze. Sequentially, we expand the dataset by resampling the continuous parameters of circuits consisting of at least one parameterized gate six times. Note that this step means we have multiple copies of the same circuits in the dataset, but each with different parameters. With this method we increased the dataset size of 11M unique circuits to 63M unitaries, which we then use to train the model presented in Section 4. Our final dataset has a memory size of $\sim 290$ GB in total. Notably, to reduce training time and storage space, we store the unitaries in half precision (float16), which introduces a mean absolute error of $\mathrm{MAE}(UU^{\dagger},I)\approx 10^{-5}$ into the unitary property $UU^{\dagger}=I$. We assume that this is the potential reason of why the infidelities presented in Fig. 3b are mostly distributed above $10^{-5}$. Future works can explore training with higher precision unitaries.
@@
The dataset creation and verification was done via CUDA-Q [cudaq], an open-source QPU-agnostic platform designed for accelerated quantum supercomputing. By offering a unified programming model in Python and C++ for co-located GPUs, QPUs, and CPUs, CUDA-Q enables the integration of classical and quantum resources within a single application, ensuring optimal performance and efficiency. The platform includes the NVQ++ compiler, which supports split compilation by lowering quantum kernels into multi-level intermediate representation (MLIR) and quantum intermediate representation (QIR). This approach ensures tight coupling between classical and quantum operations, facilitating accelerated execution of large-scale quantum workloads.
@@
Appendix D Training details
We train the Circuit-Diffusion-Transformer (CirDiT) (see App. F) from Section 4 with the dataset detailed in App. C. For this, we optimize the loss defined in Eq. 7 using the Adam optimizer  [kingma2014adam], with a learning rate of $10^{-4}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$, together with a one-cycle learning rate strategy [smith2019super]. Training is performed on 16 NVIDIA A100 GPUs for $\sim 500$k update steps, with an effective batch size of 2048. The training time is roughly 800 GPU hours. Additionally, as mentioned in Section 3.1, we drop the condition $\mathbf{c}$ with a probability of $10\%$, following classifier-free guidance (CFG) [ho2022classifier].
For the multimodal diffusion process, we use for both modes the same number of times steps $T_{h}=T_{w}=T=1000$ (see details in App. A). We use for the discrete mode a learned discrete noise schedule for the linear target (see Fig. 2ac and Section A.6) and the CosineAlpha2 schedule for the continuous mode (see Fig. 2bc and Section B.2).
During training, to reduce the variance of the ELBO estimation, we sample the timesteps $t$ and $\tilde{t}$ of Eq. 7 using a low-discrepancy sampling method, as done in Refs. [sahoo2024simple, kingma2021variational]. Considering a batch size of $m$, we sample both diffusion times for each batch sample $i$ by
@@ Appendix D Training details
As the loss in Eq. 7 requires two independent times, i.e. $t$ and $\tilde{t}$, we shuffle $\tilde{t}$ across the batch samples after sampling them. Furthermore, to increase the training time spend on similar times $t\approx\tilde{t}$, we do not shuffle the time step bins of $\tilde{t}$ with a probability of $5\%$, leaving $t$ and $\tilde{t}$ close together as they are sampled from the same bins, which means they differ by a maximal value of $1/m$.
Appendix E Inference details
We sample the results from Section 4 using the joint sampling model, as explained in Section A.1, for 40 time steps. As diffusion sampler, we use the CFG++ [chung2024cfgPP] variant of DPM++2M [lu2022dpm].
E.1 Balanced testset
In Section 4.2, we showed the compilation of random unitaries from a test set. This test set is split from the training dataset (detailed in App. C), ensuring that the test unitaries are indeed compilable with the available gate set. Importantly, we do this before starting training and before resampling the parameters of the parametrized gates (see App. C), guaranteeing a benchmark set of unitaries resulting from unique circuits the model has never seen during training. Additionally, we balance the test set such that we have an equal amount of circuits per gate count.
E.2 Multimodal CFG
As explained in Section 3.1, the reverse transitions for each of the modes (see Eq. 6) have two conditions: one related to the opposite mode, and another one accounting for the external condition $\mathbf{c}$. For sampling using CFG [ho2022classifier], we extend the guidance to two conditions.
Considering the guidance scales $\gamma_{h},\gamma_{w}$ and $\lambda_{h},\lambda_{w}$, we can write the guided velocity prediction (see Section A.4) for the discrete mode as
@@ E.2 Multimodal CFG
where ${\bf\it\epsilon}_{h},{\bf\it\epsilon}_{w}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\phi$ represents an empty condition. In practice, we implement $\phi$ as a trainable latent vector. Empirically, we found the values $\gamma_{h}=0.3$, $\gamma_{w}=0.1$, $\lambda_{h}=1.0$ and $\lambda_{w}=0.35$ work well for circuit generation, which we then use for the results presented in Section 4.
Appendix F Model architecture
In Fig. 7 we present the model architecture considered in this work, named Circuit-Diffusion-Transformer (CirDiT). Our model choice is based on the diffusion transformer (DiT) architecture [peebles2023scalable], and contains 257 million trainable parameters. For the encoder and decoder blocks, we use a channel size of 384, a number of blocks $N_{e}=N_{d}=6$ and a number of 12 heads. The core transformer has 1024 channels, $N_{c}=12$ blocks and 16 heads.
We inject the diffusion times $t$ and $\tilde{t}$ using Time-dependent Self-Attention (TMSA) [hatamizadeh2024diffit], using a $t_{\text{emb}}$ size of 512 channels. The time dimension of the circuit embeddings (see App. B) is encoded using the rotational position encoding $p$-RoPE of Ref. [barbero2025round], setting the parameter of the latter to $p=0.9$. In addition, for the qubit dimension, we add a learned position encoding before the encoder blocks. As visualized in Fig. 7, we replace the Layer-Norm layers with RMS-Norms [zhang2019root] and push them to the end of the blocks, mitigating the divergence problem of large scale DiTs, as discussed in  [pmlr-v119-xiong20b, zhuo2024lumina]. We use in some layers the adaptive version AdaRMS-Norm, where we scale the normalized output by an external signal which is constrained by the tanh activation [zhuo2024lumina].
@@ Appendix F Model architecture
Appendix G Unitary encoder pre-training
As explained in Section 3.1, in order to create an appropriate unitary conditioning, we pre-train a unitary encoder which encodes a given unitary $U$ together with a prompt into a condition $\mathbf{c}$. The main goal of this encoding is to align the representation of a given unitary to its circuit representation, ensuring that the information passed to the DM is as expressive as possible given the generation task.
Inspired by the contrastive loss of the CLIP framework [clip], we match the latent encodings of a circuit encoder and a unitary-prompt encoder, as presented in Fig. 8a. Additionally, we first encode the text prompts using a frozen pre-trained OpenCLIP [ilharcoOpenCLIP] model, specifically the architecture ViT-B-32 trained on the dataset datacomp_xl_s13b_b90k. The circuit and unitary encoder are based on the diffusion transformer architecture [peebles2023scalable]. The whole UnitaryCLIP (both encoders) we use in Section 4 contains 38 million trainable parameters. We use an additive absolute position encoding [Attention] for the unitary matrix elements.
We optimize the UnitaryCLIP using the Adam optimizer  [kingma2014adam], with a learning rate of $3.2\cdot 10^{-4}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$, together with a one-cycle learning rate strategy [smith2019super]. We train for $\sim 500$k updates steps, with an effective batch size of 4096, on the dataset of App. C. We find that the trained UnitaryCLIP is able to achieve a 99% correct matching of unseen unitary-circuit pairs.
@@ Appendix G Unitary encoder pre-training
Appendix H Additional figure parameters
We list in Table 1 additional sample parameters of the figures presented throughout this paper.
Table 1: Figure sampling parameters.
Figure Number of unitaries Circuit samples per unitary Notes
Fig. 3 2048 128 # unitaries per qubit count
Fig. 4ab - 128 A unitary per grid point
Fig. 4cd - 128 A unitary per time step
Fig. 5abc 1 2048 QFT for 4 qubits
Fig. 5defghi - - Same circuits as in Fig. 4ab for 4 qubits
Fig. 9 16384 1 Here # unitaries is # ciruits
Appendix I Circuit corruption test
In Fig. 3a we observed the appearance of some characteristic peaks in the infidelity distribution. In order to further investigate the origin of such peaks, we perform different corruptions on test set circuits and record the resulting infidelities between corrupted and original circuits. We show in Fig. 9 the infidelities for the following discrete and continuous corruptions:
@@ Appendix I Circuit corruption test
Comparing the peaks in the discrete corruptions of Fig. 9a to the ones observed in the random unitary compilation (see Section 4.2), we see the same peaks appearing at $0.4$ and $0.8$. Hence, we attribute the peaks arising in Fig. 3a to cases in which the model incorrectly places gates in a similar way as our corruption tests, this means, by misplacing a single gate. On the other hand, corrupting the continuous values of the parameterized gates causes a wide continuous distribution around zero (Fig. 9b), where the peak at zero infidelity corresponds to circuits which have no parameterized gates. This highlights an important result: comparing this distribution to that in Fig. 3a showcases the fact that the model predicts the continouous parameters rather accurately, as no broad distribution arises, and the error arise mainly from misplacements of a single gate.
Figure 9: Circuit corruptions. a) Infidelity distribution for circuits with discrete corruptions (see App. I). b) Infidelity distribution for circuits with continuous corruption, as defined in Eq. 70.
Appendix J Hamiltonians
In Section 4.3 we showed the compilation of the evolution operators of the Ising and XXZ Hamiltonians. Here we define these Hamiltonians in terms of the Pauli operators, which are defined as
@@
where $J\in\mathbb{R}$ is the coupling constant, $\Delta\in\mathbb{R}$ a perturbation and $h\in\mathbb{R}$ a magnetic field. In Section 4, we fix $h=0.2$ for the XXZ Hamitonian.
Appendix K Additional GPE structures
In this section, we present additional circuit structures extracted using our proposed Gate-Pair Encoding (GPE) scheme (see Section 3.4), extending the results presented in Section 4.4. We follow the exact same recipe as in the main text, and take all circuits generated by the DM, without any selection or filtering for wrong circuits. Notably, this means all structures are extracted without any circuit evaluation, keeping all computations of the process purely classic.
For the structures shown in Fig. 10, Fig. 11 and Fig. 12, we run GPE for a maximum of 250 iterations, or until there are no pairs left to be matched into a new token. We present structures for different token depths, where depth 0 is defined as the elementary tokens, being here the original gate set (e.g. $\mathrm{h}$ or $\mathrm{cx}$). Then, depth 1 structures are gate-pairs (i.e. pairs of depth 0 tokens). Further, depth 2 tokens are constructed from at least one depth 1 token together with either one depth 1 or 0 token. More generally, a depth $m$ token always consists of one depth $m-1$ token and another one with depth $\leq m-1$.
Figure 10: QFT Gate-Pair Encodings. GPE (see Section 3.4) on generated circuits for the QFT unitary for a) 3 qubits, b) 4 qubits and c) 5 qubits. The rows correspond to the token depth (see App. K). The circuits shown represent the top-5 most occurring structures, from most occurring (left) to less often occurring (right).
Figure 11: Ising Hamiltonian Gate-Pair Encodings. GPE (see Section 3.4) on generated circuits for the 4-qubit Ising Hamiltonian (defined in Eq. 72) evolution unitary for $\tau=0.25$. We use in a) circuits for the parameters $h\in\left[0.5,0.9\right]$ and $J=0$, and in b) $J\in\left[0.5,0.9\right]$ and $h=0$. The rows correspond to the token depth (see App. K). The circuits shown represent the top-5 most occurring structures, from most occurring (left) to less often occurring (right).
Figure 12: XXZ Hamiltonian Gate-Pair Encodings. GPE (see Section 3.4) on generated circuits for the 4-qubit XXZ Hamiltonian (defined in Eq. 73) evolution unitary for $\tau=0.25$. We use in a) circuits for the parameters $J\in\left[0.5,0.9\right]$ and $\Delta=0$, and in b) $J\in\left[0.5,0.9\right]$ and $\Delta\in\left[0.5,0.9\right]$. The rows correspond to the token depth (see App. K). The circuits shown represent the top-5 most occurring structures, from most occurring (left) to less often occurring (right).

### v2

@@ Appendix D Dataset details
In this section, we present the details of the training dataset used to train the model in Section˜IV.
@@
We create the dataset by generating random circuits consisting of the gates $\left\{\mathrm{h},\mathrm{cx},\mathrm{ccx},\mathrm{swap},\mathrm{rx},\mathrm{ry},\mathrm{rz},\mathrm{cp}\right\}$ with CUDA-Q [48] (see details on the latter below). For this, we first sample a subset of the gates from the previous set, and then append these gates uniformly to an empty circuit. We do this randomly from 4 to 32 gates and for 3 to 5 qubits. Once the gate type is fixed, we sample the continuous parameters, of the parameterized gates, following a uniform distribution. After a circuit is sampled, we evaluate the implemented unitary and store the circuit-unitary pair. Additionally, we parse the sampled gate subset into a text prompt and store it alongside the circuits.
We sample random circuits until our dataset consists of 11 million (M) unique circuit ansätze. Sequentially, we expand the dataset by resampling the continuous parameters of circuits consisting of at least one parameterized gate six times. Note that this step means we have multiple copies of the same circuits in the dataset, but each with different parameters. With this method we increased the dataset size of 11M unique circuits to 63M unitaries, which we then use to train the model presented in Section˜IV. Our final dataset has a memory size of $\sim$290 GB in total. Notably, to reduce training time and storage space, we store the unitaries in half precision (float16), which introduces a mean absolute error of $\mathrm{MAE}(UU^{\dagger},I)\approx 10^{-5}$ into the unitary property $UU^{\dagger}=I$. We assume that this is the potential reason of why the infidelities presented in Fig.˜3b are mostly distributed above $10^{-5}$. Future works can explore training with higher precision unitaries.
@@
The dataset creation and verification was done via CUDA-Q [48], an open-source QPU-agnostic platform designed for accelerated quantum supercomputing. By offering a unified programming model in Python and C++ for co-located GPUs, QPUs, and CPUs, CUDA-Q enables the integration of classical and quantum resources within a single application, ensuring optimal performance and efficiency. The platform includes the NVQ++ compiler, which supports split compilation by lowering quantum kernels into multi-level intermediate representation (MLIR) and quantum intermediate representation (QIR). This approach ensures tight coupling between classical and quantum operations, facilitating accelerated execution of large-scale quantum workloads.
@@
Appendix E Training details
We train the Circuit-Diffusion-Transformer (CirDiT) (see App.˜G) from Section˜IV with the dataset detailed in App.˜D. For this, we optimize the loss defined in Eq.˜4 using the Adam optimizer  [50], with a learning rate of $10^{-4}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$, together with a one-cycle learning rate strategy [51]. Training is performed on 16 NVIDIA A100 GPUs for $\sim$800k update steps, with an effective batch size of 2048. The training time is roughly 700 single GPU hours. Additionally, as mentioned in Section˜III.1, we drop the condition $\mathbf{c}$ with a probability of $10\%$, following classifier-free guidance (CFG) [69].
For the multimodal diffusion process, we use for both modes the same number of times steps $T_{h}=T_{w}=T=1000$ (see details in App.˜A). We use for the discrete mode a learned discrete noise schedule for the linear target (see Fig.˜2ac and Section˜A.6) and the cosine squared schedule for the continuous mode (see Fig.˜2bc and Section˜B.2).
During training, to reduce the variance of the ELBO estimation, we sample the timesteps $t$ and $\tilde{t}$ of Eq.˜4 using a low-discrepancy sampling method, as done in Refs. [65, 68]. Considering a batch size of $m$, we sample both diffusion times for each batch sample $i$ by
@@ Appendix E Training details
As the loss in Eq.˜4 requires two independent times, i.e. $t$ and $\tilde{t}$, we shuffle $\tilde{t}$ across the batch samples after sampling them. Furthermore, to increase the training time spend on similar times $t\approx\tilde{t}$, we do not shuffle the time step bins of $\tilde{t}$ with a probability of $5\%$, leaving $t$ and $\tilde{t}$ close together as they are sampled from the same bins, which means they differ by a maximal value of $1/m$.
Appendix F Inference details
We sample the results from Section˜IV using the joint sampling model, as explained in Section˜A.1, for 40 time steps. As diffusion sampler, we use the CFG++ [52] variant of DPM++2M [53].
The benchmark runtime values, reported in Table˜1, are computed on a workstation with the AMD Ryzen Threadripper PRO 7955WX CPU and a NVIDIA RTX PRO 6000 Blackwell Max-Q GPU.
F.1 Balanced testset
In Section˜IV.2, we showed the compilation of random unitaries from a test set. This test set is split from the training dataset (detailed in App.˜D), ensuring that the test unitaries are indeed compilable with the available gate set. Importantly, we do this before starting training and before resampling the parameters of the parametrized gates (see App.˜D), guaranteeing a benchmark set of unitaries resulting from unique circuits the model has never seen during training. Additionally, we balance the test set such that we have an equal amount of circuits per gate count.
F.2 Multimodal CFG
As explained in Section˜III.1, the reverse transitions for each of the modes (see Eq.˜13) have two conditions: one related to the opposite mode, and another one accounting for the external condition $\mathbf{c}$. For sampling using CFG [69], we extend the guidance to two conditions.
Considering the guidance scales $\gamma_{h},\gamma_{w}$ and $\lambda_{h},\lambda_{w}$, we can write the guided velocity prediction (see Section˜A.4) for the discrete mode as
@@ F.2 Multimodal CFG
where ${\bf\it\epsilon}_{h},{\bf\it\epsilon}_{w}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\phi$ represents an empty condition. In practice, we implement $\phi$ as a trainable latent vector. Empirically, we found the values $\gamma_{h}=0.3$, $\gamma_{w}=0.1$, $\lambda_{h}=1.0$ and $\lambda_{w}=0.35$ work well for circuit generation, which we then use for the results presented in Section˜IV.
Appendix G Model architecture
@@
In Fig.˜14 we present the model architecture considered in this work, named Circuit-Diffusion-Transformer (CirDiT). Our model choice is based on the diffusion transformer (DiT) architecture [49], and contains 151 million trainable parameters. For the encoder and decoder blocks, we use a channel size of 256, a number of blocks $N_{e}=N_{d}=6$ and a number of 8 heads. The core transformer has 768 channels, $N_{c}=12$ blocks and 16 heads.
We inject the diffusion times $t$ and $\tilde{t}$ using Time-dependent Self-Attention (TMSA) [74], using a $t_{\text{emb}}$ size of 512 channels. The time dimension of the circuit embeddings (see App.˜B) is encoded using the rotational position encoding $p$-RoPE of Ref. [75], setting the parameter of the latter to $p=0.9$. In addition, for the qubit dimension, we add a learned position encoding before the encoder blocks. As visualized in Fig.˜14, we replace the Layer-Norm layers with RMS-Norms [76] and push them to the end of the blocks, mitigating the divergence problem of large scale DiTs, as discussed in  [77, 78]. We use in some layers the adaptive version AdaRMS-Norm, where we scale the normalized output by an external signal which is constrained by a $\mathrm{tanh}$ activation [78].
Appendix H Unitary encoder pre-training
As explained in Section˜III.1, in order to create an appropriate unitary conditioning, we pre-train a unitary encoder which encodes a given unitary $U$ together with a prompt into a condition $\mathbf{c}$. The main goal of this encoding is to align the representation of a given unitary to its circuit representation, ensuring that the information passed to the DM is as expressive as possible given the generation task.
Inspired by the contrastive loss of the CLIP framework [79], we match the latent encodings of a circuit encoder and a unitary-prompt encoder, as presented in Fig.˜15a. Additionally, we first encode the text prompts using a frozen pre-trained OpenCLIP [45] model, specifically the architecture ViT-B-32 trained on the dataset datacomp_xl_s13b_b90k. The circuit and unitary encoder are based on the diffusion transformer architecture [49]. The whole UnitaryCLIP (both encoders) we use in Section˜IV contains 38 million trainable parameters. We use an additive absolute position encoding [80] for the unitary matrix elements.
We optimize the UnitaryCLIP using the Adam optimizer  [50], with a learning rate of $3.2\cdot 10^{-4}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$, together with a one-cycle learning rate strategy [51]. We train for $\sim 550$k updates steps, with an effective batch size of 4096, on the dataset of App.˜D. We find that the trained UnitaryCLIP is able to achieve a 99% correct matching of unseen unitary-circuit pairs.
@@ Appendix H Unitary encoder pre-training
Appendix I Additional figure parameters
We list in Table˜7 additional sample parameters of the figures presented throughout this paper.
Table 7: Additional figure and table sampling parameters.
Number of unitaries Circuit samples per unitary Notes
Fig.˜3ab 1024 128 # unitaries per qubit count
Fig.˜3c 2048 128 # unitaries per qubit count
Fig.˜5 480 1 # unitaries per qubit count
Fig.˜6 480 64 # unitaries per qubit count
Fig.˜7ab - 128 A unitary per grid point
Fig.˜7cd - 128 A unitary per time step
Fig.˜8abc 1 2048 QFT for 4 qubits
Fig.˜8defghi - - Same circuits as in Fig.˜7ab for 4 qubits
Fig.˜12 - - Same data as Fig.˜5
Fig.˜16 16384 1 Here # unitaries is # ciruits
Table˜1 480 128 # unitaries per qubit count
Table˜4 480 64 # unitaries per qubit count
Table˜5 480 64 # unitaries per qubit count
Appendix J Circuit corruption test
In Fig.˜3a we observed the appearance of some characteristic peaks in the infidelity distribution. In order to further investigate the origin of such peaks, we perform different corruptions on test set circuits and record the resulting infidelities between corrupted and original circuits. We show in Fig.˜16 the infidelities for the following discrete and continuous corruptions:
@@ Appendix I Additional figure parameters
Comparing the peaks in the discrete corruptions of Fig.˜16a to the ones observed in the random unitary compilation (see Section˜IV.2), we see the same peaks appearing at $0.4$ and $0.8$. Hence, we attribute the peaks arising in Fig.˜3a to cases in which the model incorrectly places gates in a similar way as our corruption tests, this means, by misplacing a single gate. On the other hand, corrupting the continuous values of the parameterized gates causes a wide continuous distribution around zero (Fig.˜16b), where the peak at zero infidelity corresponds to circuits which have no parameterized gates. This highlights an important result: comparing this distribution to that in Fig.˜3a showcases the fact that the model predicts the continouous parameters rather accurately, as no broad distribution arises, and the error arise mainly from misplacements of a single gate.
Figure 16: Circuit corruptions. a) Infidelity distribution for circuits with discrete corruptions (see App.˜J). b) Infidelity distribution for circuits with continuous corruption, as defined in Eq.˜71.
Appendix K Noisy simulation
To test the noise robustness of the synthesized circuits, we simulate them under the influence of gate errors. We then calculate the process infidelity as a measure to estimate how well the circuits could run on actual NISQ hardware, as we show in Fig.˜4.
In the noiseless simulation, we can write the channel of a target unitary $U$ as
$\mathcal{E}_{U}(\rho)=U\rho U^{\dagger},$ (72)
where $\rho$ is the density matrix. Next, we define the process infidelity–a distance between the target channel $\mathcal{E}_{U}$ and the noisy channel $\mathcal{E}_{\text{noisy}}$, which is given by a synthesized circuit in the presence of gate errors–as
$\mathcal{I}(\mathcal{E}_{\text{noisy}},\mathcal{E}_{U})=1-\frac{1}{d^{2}}\mathrm{Tr}\left[S_{\mathcal{E}_{U}}^{\dagger}\mathcal{S}_{\mathcal{E}_{\text{noisy}}}\right],$ (73)
where $d$ is the dimension of the channels. Here, $S_{\mathcal{E}_{U}}$ and $\mathcal{S}_{\mathcal{E}_{\text{noisy}}}$ are the super-operator representations of the channels.
As a simplified model, we describe the noisy circuits with depolarizing errors applied to all gates. Specifically, we assign each $k$-qubit gate, with unitary $U_{\text{gate}}=U_{k}\otimes\mathbf{I}$ and connections on the qubits $\{k_{i}\}$, placed in a $n$-qubit circuit, a depolarizing error given by
$\mathcal{E}_{k}(\rho)=(1-p_{k})\;U_{\text{gate}}\rho U_{\text{gate}}^{\dagger}+p_{k}\;\frac{1}{2^{k}}\mathbf{I}_{\{k_{i}\}}\otimes\mathrm{Tr}_{\{k_{i}\}}[\rho],$ (74)
where the error acts simultaneously on the $k$ qubits. Here, $\mathrm{Tr}_{\{k_{i}\}}$ is the trace over all $k$ qubits the gate acts on, $\mathbf{I}_{\{k_{i}\}}$ is the corresponding identity matrix, and $p_{k}$ is the probability that the depolarizing error occurs.
To parametrize the error probabilities $p_{k}$ with a single control parameter $p$, we model the probability that a $k$-qubit gate acts without error as the probability that $k$ independent single qubit gates act without an error, i.e. we set
$1-p_{k}\overset{!}{=}(1-p_{1})^{k}.$ (75)
This relation then results in the equations:
$p_{1}=p,\quad p_{2}=1-(1-p)^{2}\text{ and}\quad p_{3}=1-(1-p)^{3},$ (76)
which we use for the results shown in Fig.˜4.
Appendix L Hamiltonians
In Section˜IV.3 we showed the compilation of the evolution operators of the Ising and XXZ Hamiltonians. Here we define these Hamiltonians in terms of the Pauli operators, which are defined as
@@
where $J\in\mathbb{R}$ is the coupling constant, $\Delta\in\mathbb{R}$ a perturbation and $h\in\mathbb{R}$ a magnetic field. In Section˜IV, we fix $h=0.2$ for the XXZ Hamitonian.
Appendix M Additional GPE structures
In this section, we present additional circuit structures extracted using our proposed Gate-Pair Encoding (GPE) scheme (see Section˜III.4), extending the results presented in Fig.˜8. We follow the exact same recipe as in the main text, and take all circuits generated by the DM, without any selection or filtering for wrong circuits. Notably, this means all structures are extracted without any circuit evaluation, keeping all computations of the process purely classic.
For the structures shown in Fig.˜17, Fig.˜18, Fig.˜19 and Fig.˜20, we run GPE for a maximum of 250 iterations, or until there are no pairs left to be matched into a new token. We present structures for different token depths, where depth 0 is defined as the elementary tokens, being here the original gate set (e.g. $\mathrm{h}$ or $\mathrm{cx}$). Then, depth 1 structures are gate-pairs (i.e. pairs of depth 0 tokens). Further, depth 2 tokens are constructed from at least one depth 1 token together with either one depth 1 or 0 token. More generally, a depth $m$ token always consists of one depth m1m-1 token and another one with depth $\leq m-1$.
We present in Fig.˜17, Fig.˜18 and Fig.˜19 structures generated by our method (genqc2), compared to structures from other methods in Fig.˜20.
Figure 17: QFT Gate-Pair Encodings of genqc2. GPE (see Section˜III.4) on generated circuits for the QFT unitary for a) 3 qubits, b) 4 qubits and c) 5 qubits. The rows correspond to the token depth (see App.˜M). The circuits shown represent the top-5 most occurring structures, from most occurring (left) to less often occurring (right).
Figure 18: Ising Hamiltonian Gate-Pair Encodings of genqc2. GPE (see Section˜III.4) on generated circuits for the 4-qubit Ising Hamiltonian (defined in Eq.˜78) evolution unitary for $\tau=0.25$. We use in a) circuits for the parameters $h\in\left[0.5,0.9\right]$ and $J=0$, and in b) $J\in\left[0.5,0.9\right]$ and $h=0$. The rows correspond to the token depth (see App.˜M). The circuits shown represent the top-5 most occurring structures, from most occurring (left) to less often occurring (right).
Figure 19: XXZ Hamiltonian Gate-Pair Encodings of genqc2. GPE (see Section˜III.4) on generated circuits for the 4-qubit XXZ Hamiltonian (defined in Eq.˜79) evolution unitary for $\tau=0.25$. We use in a) circuits for the parameters $J\in\left[0.5,0.9\right]$ and $\Delta=0$, and in b) $J\in\left[0.5,0.9\right]$ and $\Delta\in\left[0.5,0.9\right]$. The rows correspond to the token depth (see App.˜M). The circuits shown represent the top-5 most occurring structures, from most occurring (left) to less often occurring (right).
Figure 20: QFT Gate-Pair Encodings. GPE (see Section˜III.4) on generated circuits for the 4-qubit QFT unitary for a) QSD, b) AQC, c) LEAP and d) QFAST. The rows correspond to the token depth (see App.˜M). The circuits shown represent the top-3 most occurring structures, from most occurring (left) to less often occurring (right).